OpenStudy (anonymous):

find percent notation of 6.2 simp your answer type interger or decimal

OpenStudy (anonymous):

62/10 = 620/100 = 620%

OpenStudy (saifoo.khan):

620%

OpenStudy (anonymous):

It is already in decimal form

OpenStudy (anonymous):

oh i just wrote everything it said

OpenStudy (anonymous):

so it is 620%

OpenStudy (saifoo.khan):

it says percentage notation form.

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

It must be 620%

OpenStudy (anonymous):

ok thanks

OpenStudy (saifoo.khan):

uw.

OpenStudy (anonymous):

welcome